### v2

# ScaleOpt: GPU-Accelerated and Scalable Optimization

Abstract
We propose a reduced-space formulation for optimizing over trained neural networks where the network’s outputs and derivatives are evaluated on a GPU. To do this, we treat the neural network as a “gray box” where intermediate variables and constraints are not exposed to the optimization solver. Compared to the full-space formulation, in which intermediate variables and constraints are exposed to the optimization solver, the reduced-space formulation leads to faster solves and fewer iterations in an interior point method. We demonstrate the benefits of this method on two optimization problems: Adversarial generation for a classifier trained on MNIST images and security-constrained optimal power flow with transient feasibility enforced using a neural network surrogate.
1 Introduction
Optimization over trained machine learning (ML) models can be used to verify ML models [tjeng2018evaluating, bunel2018unified], generate adversarial examples, and use these ML models for optimization-based design and control [bugosen2024, lopezflores2024]. For smooth, nonlinear ML models (e.g., neural networks with non-ReLU activation functions) or design or control tasks involving nonconvex constraints, the resulting optimization problem must be solved with a nonconvex optimization solver (as opposed to a mixed-integer, linear solver). Global optimization of nonconvex functions is well-known to be NP-hard, but local optimization of these functions can be done efficiently with interior point methods [nocedalwright].
Despite favorable scalability of interior point methods in other domains (e.g., power grid operation [aravena2023]), we are unaware of recent work investigating the scalability of interior point methods with large neural networks (NNs) embedded. Recent work using interior point methods to optimize over trained NNs [kilwein2023, bugosen2024, casas2025comparison] has been limited to small NN models (fewer than 1M trained parameters). In this work, we examine the scalability of interior point methods with large NN models (over 100M trained parameters) embedded. We propose a reduced-space formulation with GPU acceleration as a scalable method of solving these optimization problems.
GPU-accelerated optimization has received much recent attention. First-order methods for linear programming [applegate2021pdhg] have led to implementations that offload matrix-vector multiplications to a GPU [lu2024cupdlp]. Simultaneously, the development of GPU-accelerated sparse matrix factorization algorithms, e.g., cuDSS [cudss], have enabled GPU-acceleration for a wider class of optimization problem [shin2024accelerating, pacaud2024gpu]. In contrast to these general-purpose methods, our method is tailored to nonlinear optimization problems with neural networks embedded.
2 Nonlinear optimization
We study nonlinear optimization problems in the form given by Problem (1):
$\min_{x}f(x)\text{ subject to }\left\{\begin{array}[]{l}g(x)=0\\ x\geq 0\end{array}\right$. (1)
This formulation does not preclude general inequality constrains, which may be reformulated with a slack variable. We consider interior point methods, such as IPOPT [ipopt], for solving (1), which require that functions $f$ and $g$ are twice continuously differentiable [nocedalwright]. We note that this requirement precludes neural networks with non-differentiable activation functions such as ReLU. Interior point methods iteratively compute search directions $d$ by solving the linear system (2):
$\begin{bmatrix}(\nabla^{2}\mathcal{L}(x)+\alpha)&\nabla g(x)^{T}\\ \nabla g(x)&0\end{bmatrix}d=-\begin{bmatrix}\nabla f(x)+\nabla g(x)^{T}\lambda+\beta\\ g(x)\end{bmatrix}$, (2)
where $\alpha$ and $\beta$ are additional terms that are not shown for simplicity. The matrix on the left-hand side is referred to as the Karush-Kuhn Tucker, or KKT, matrix. To construct this system, solvers rely on callbacks (oracles) that provide the Jacobian, $\nabla g$, and the Hessian of the Lagrangian function, $\nabla^{2}\mathcal{L}$. These are typically provided by an automatic differentiation system [griewank2008].
3 Representing neural networks as constraints
We consider neural network predictors, $y={\rm NN}(x)$, defined by repeated application of an affine transformation and a nonlinear activation function $\sigma$ over $L$ layers:
$y_{l}$ $=\sigma_{l}(W_{l}y_{l-1}+b_{l})$ $l\in\{1,\dots,L\},$ (3)
where $y_{0}=x$ and $y=y_{L}$. In this context, training weights $W_{l}$ and $b_{l}$ are considered constant. Pre-trained neural network predictors may be embedded into the constraints of an optimization problem using either full-space or reduced-space formulations [omlt, schweidtmann2019deterministic].
3.1 Full-space
In the full-space formulation, we add additional variables and constraints for each layer of the neural network:
$z_{l}$ $={W_{l}}y_{l-1}+b_{l}$ $l\in\{1,\dots,L\}$ (4)
$y_{l}$ $=\sigma_{l}(z_{l})$ $l\in\{1,\dots,L\}.$
The full-space approach prioritizes small, sparse nonlinear constraints at the cost of introducing (potentially) many new variables and constraints.
3.2 Reduced-space
In the reduced-space formulation, we add a single vector-valued decision variable $y$ to represent the output of the final activation function and a single vector-valued nonlinear equality constraint that encodes the complete neural network:
$y={\rm NN}(x)=\sigma_{L}({W_{L}}(\ldots\sigma_{l}({W_{l}}(\ldots\sigma_{1}({W}_{1}x+b_{1})\ldots)+b_{l})\ldots)+b_{L})$ (5)
The benefit of this formulation is that we only add a single decision variable and a single nonlinear equality constraint (each of dimension of the neural network’s output). The drawback is that this nonlinear equality constraint contains a large, complicated algebraic expression that can be expensive to evaluate and differentiate. However, this can be alleviated by using dedicated neural network modeling software (e.g., PyTorch [paszke2019pytorch]) to represent the neural network constraint rather than general-purpose algebraic modeling software (e.g., JuMP [jump1]). Doing so also allows us to exploit GPU acceleration via PyTorch’s CUDA interface. We note that this approach limits the information that can be communicated between the neural network and the optimization solver. Oracles (which PyTorch provides for function, Jacobian, and Hessian evaluation) can be queried, but the internal structure of the neural network cannot be exploited directly. This is suitable for nonlinear local optimization, where only oracles are required, but not for global optimization, where the algebraic form of constraints is exploited to construct relaxations.
3.2.1 The Hessian of the Lagrangian
Interior point methods require oracles to evaluate the Hessian of the Lagrangian, $\nabla^{2}\mathcal{L}$:
$\nabla^{2}\mathcal{L}(x,\lambda)=\nabla^{2}f(x)+\sum_{i}\lambda_{i}\nabla^{2}g_{i}(x),$ (6)
where $\lambda$ is the vector of Lagrange multipliers of the equality constraints in Equation 1. Naively constructing $\sum_{i}\lambda_{i}\nabla^{2}g_{i}$ for constraints involving a reduced-space neural network, $g(x,y)=y-{\rm NN}(x)$, would require (1) evaluating the Hessian of the neural network, $\nabla^{2}{\rm NN}(x)$, and (2) computing a sum-product with $\lambda$ along the first rank of this third-order Hessian tensor. This sum-product is potentially expensive for a dense Hessian ($\mathcal{O}(mn^{2})$, where $m$ is the output dimension and $n$ is the input dimension of the neural network). For this reason, we encode the Lagrangian of the neural network, $\lambda^{T}{\rm NN}(x)$, directly as a linear layer in PyTorch and differentiate through this scalar-valued function to directly compute the $n\times n$ Hessian matrix, $\sum_{i}\lambda_{i}\nabla^{2}{\rm NN}_{i}(x)$.
4 Test problems
We test the full and reduced-space formulations on two nonlinear optimization problems with NNs embedded: (1) Adversarial image generation using a neural network MNIST classifier (denoted “MNIST”) and (2) security-constrained optimal power flow with a neural network constraint enforcing transient feasibility (denoted “SCOPF”).
Table 1: Structure of neural network models
Model N. inputs N. outputs N. neurons N. param. Activation
MNIST 784 10 1k 167k Tanh+SoftMax
MNIST 784 10 3k 1M Tanh+SoftMax
MNIST 784 10 5k 5M Tanh+SoftMax
MNIST 784 10 11k 18M Tanh+SoftMax
MNIST 784 10 21k 70M Tanh+SoftMax
MNIST 784 10 41k 274M Sigmoid+SoftMax
SCOPF 117 37 254 15k Tanh
SCOPF 117 37 1k 578k Tanh
SCOPF 117 37 5k 4M Tanh
SCOPF 117 37 36k 68M Tanh
SCOPF 117 37 152k 592M Tanh
4.1 Adversarial image generation for an MNIST classifier
We train a set of neural networks using smooth activation functions (hyperbolic tangent, sigmoid, and softmax) to serve as classifiers for images from the MNIST set of handwritten digits [lecun1998mnist]. Inputs are the $\times28$ grayscale pixel colors, flattened into a 784-dimensional vector, and outputs are scores for each digit, 0-9, that may be interpreted as the probability that the image represents the corresponding digit. The neural networks each have seven layers total and have between 128 and 8192 neurons per hidden layer. The networks are trained to have accuracies of at least 95% on the test set of 10,000 images from the dataset. The number of neurons, trained parameters, and activation functions for each network are shown in Table 1. Our optimization problem finds a minimal perturbation to a reference image that results in a misclassification:
[…]
Here, $x$ contains the grayscale values of the generated image, $x_{\rm ref}$ contains those of the reference image, and $t$ is the coordinate of the neural network’s output, $y$, corresponding to a target label (misclassification). The neural network constraint, $y={\rm NN}(x)$, may be written in full-space or reduced-space formulations. The image must be misclassified with at least 60% confidence.
4.2 Transient-constrained optimal power flow
Security-constrained optimal power flow (SCOPF) is a well-established problem for dispatching generators in an electric power network in which feasibility of the network (i.e., the ability to meet demand) is enforced for a set of contingencies [aravena2023]. Each contingency $k$ represents the loss of a set of generators and/or power lines. We consider a variant of this problem where, in addition to enforcing steady-state feasibility, we enforce feasibility of the transient response resulting from the contingency. In particular, we enforce that the transient frequency at each bus is at least ${\bf\eta}=59.4$ Hz for the 30 second interval following each contingency. This problem is given by Equation 8:
[…]
Here, $S^{g}$ is a vector of complex AC power generations for each generator in the network, $V$ is a vector of complex bus voltages, $c$ is a quadratic cost function, and $\bf S^{d}$ is a constant vector of complex power demands. Here $F_{k}\leq 0$ represents the set of constraints enforcing feasibility of the power network for contingency $k$ (see [cain2012history]), where $k=0$ refers to the base network, and $G_{k}$ maps generations and demands to the minimum frequency at each bus over the interval considered.
In this work, we consider an instance of Problem 8 defined on a 37-bus synthetic test grid [birchfield2017, hawaii40]. In this case, $G_{k}$ has 117 inputs and 37 outputs. We consider a single contingency that outages generator 5 on bus 23. We choose a small grid model with a single contingency because our goal is to test the different neural network formulations, not the SCOPF formulation itself.
Stability surrogate model Instead of considering the differential equations describing transient behavior of the power network directly in the optimization problem, we approximate $G_{k}$ with a neural network surrogate, as proposed by [garcia2025transient]. Our surrogate is trained on data from 110 high-fidelity simulations using PowerWorld [powerworldmanual] with generations and loads uniformly sampled from within a $\pm 20\%$ interval of each nominal value. We use sequential neural networks with $\tanh$ activation functions with between two and 20 layers and between 50 and 4,000 neurons per layer. As shown in Table 1, these networks have between 7,000 and 592 million trained parameters. The networks are trained to minimize mean squared error using the Adam optimizer [adam] until training loss (mean squared prediction error) is below 0.01 for 1,000 consecutive epochs. We use a simple training procedure and small amount of data because our goal is to test optimization formulations with embedded neural networks, rather than the neural networks themselves.
5 Results
5.1 Computational setting
We model the SCOPF problem using PowerModels [powermodels], PowerModelsSecurityConstrained [powermodelssecurity], and JuMP [jump1]. Neural networks are modeled using PyTorch [paszke2019pytorch] and embedded into the optimization problem using MathOptAI.jl [dowson2025moai]. Optimization problems are solved using the IPOPT interior point method [ipopt] with MA57 [ma57] as the linear solver to a tolerance of $10^{-6}$. We note that IPOPT runs exclusively on CPU. Interfacing our function evaluation methods with a GPU-enabled nonlinear optimization solver, e.g., MadNLP [shin2024accelerating], would be valuable future work. We evaluate and differentiate the full-space model’s functions on a CPU using JuMP. While we could use a framework that supports GPU-accelerated function evaluations such as PyTorch or ExaModels (see [shin2024accelerating]), our results indicate that these function evaluations are not a significant contributor to solve time (see Table 4). Because our reduced-space models use PyTorch, they can be evaluated on a CPU or GPU. We run our experiments on the Darwin cluster at Los Alamos National Laboratory. The CPU used in these experiments is an AMD EPYC with 128 cores and 500 GB of RAM and the GPU is an NVIDIA A100 with 40 GB of on-device memory. The code used to produce these results can be found at https://github.com/Robbybp/moai-examples.
5.2 Structural results
Table 2 shows the numbers of variables, constraints, and nonzero entries of the derivative matrices for the two optimization problems with different neural networks and formulations. With the reduced-space formulations, the structure of the optimization problem does not change as the neural network surrogate adds more interior layers. By contrast, the full-space formulation grows in numbers of variables, constraints, and nonzeros as the neural network gets larger. In this case, the number of nonzeros in the Jacobian matrix is approximately the number of trained parameters in the embedded neural network model. These problem structures suggest that the full-space formulation will lead to expensive KKT matrix factorizations, while this will not be an issue for the reduced-space formulation.
Table 2: Structure of optimization problems with embedded neural networks
Model Formulation NN param. N. var. N. con. Jac. NNZ Hess. NNZ
MNIST Full-space 167k 3k 2k 171k 661
MNIST Full-space 1M 7k 5k 1M 2k
MNIST Full-space 5M 12k 11k 5M 5k
MNIST Full-space 18M 22k 21k 18M 10k
MNIST Full-space 70M 43k 41k 70M 20k
MNIST Full-space 274M 84k 82k 275M 40k
MNIST Reduced-space All NNs 2k 805 10k 307k
SCOPF Full-space 15k 1k 1k 17k 1k
SCOPF Full-space 578k 4k 4k 567k 2k
SCOPF Full-space 4M 11k 11k 4M 6k
SCOPF Full-space 68M 73k 73k 68M 37k
SCOPF Full-space 592M 305k 305k 592M 153k
SCOPF Reduced-space All NNs 1k 1k 10k 8k
Table 3: Solve times with different neural networks and formulations
Model Formulation Platform NN param. Solve time (s) N. iter. Time/iter. (s) Objective
MNIST Full-space CPU 167k 19 290 0.07 3.3
MNIST Full-space CPU 1M 348 1157 0.3 3.3
MNIST Full-space CPU 5M 11536 2110 5 6.1
MNIST Full-space CPU 18M
MNIST Reduced-space CPU 167k 3 28 0.1 3.3
MNIST Reduced-space CPU 1M 4 43 0.1 3.3
MNIST Reduced-space CPU 5M 10 77 0.1 5.9
MNIST Reduced-space CPU 18M 63 147 0.4 4.7
MNIST Reduced-space CPU 70M 58 80 0.7 2.0
MNIST Reduced-space CPU 274M 45 31 1 5.3
MNIST Reduced-space CPU+GPU 167k 2 28 0.07 3.3
MNIST Reduced-space CPU+GPU 1M 4 43 0.08 3.3
MNIST Reduced-space CPU+GPU 5M 5 67 0.08 5.9
MNIST Reduced-space CPU+GPU 18M 12 149 0.08 4.8
MNIST Reduced-space CPU+GPU 70M 8 82 0.1 2.0
MNIST Reduced-space CPU+GPU 274M 3 28 0.1 5.3
SCOPF Full-space CPU 15k 4 675 0.01 82379
SCOPF Full-space CPU 578k 399 1243 0.3 82379
SCOPF Full-space CPU 4M 8858 2505 4 82379
SCOPF Full-space CPU 68M
SCOPF Reduced-space CPU 15k 7 358 0.02 82379
SCOPF Reduced-space CPU 578k 5 147 0.03 82379
SCOPF Reduced-space CPU 4M 3 53 0.06 82379
SCOPF Reduced-space CPU 68M 37 40 1 82379
SCOPF Reduced-space CPU 592M 144 42 3 82379
SCOPF Reduced-space CPU+GPU 15k 5 298 0.02 82379
SCOPF Reduced-space CPU+GPU 578k 3 162 0.02 82379
SCOPF Reduced-space CPU+GPU 4M 1 53 0.03 82379
SCOPF Reduced-space CPU+GPU 68M 2 40 0.04 82379
SCOPF Reduced-space CPU+GPU 592M 3 42 0.08 82379
Fails to solve within 5 hour time limit
5.3 Runtime results
Runtimes for the different formulations with neural network surrogates of increasing size are given in Table 3. For the reduced-space formulation, we also compare CPU-only solves with solves that leverage a GPU for neural network evaluation (and differentiation). The results immediately show that the full-space formulation is not scalable to neural networks with more than a few million trained parameters. The full-space formulation exceeds the 5-hour time limit on networks above this size. A breakdown of solve times, given in Table 4, confirms the bottleneck in this formulation. The full-space formulation spends almost all of its solve time in the IPOPT algorithm, which we assume is dominated by KKT matrix factorization.  [^1] This is difficult to confirm directly, but by parsing IPOPT’s logs, we see that, for the SCOPF model with the 4M-parameter neural network, IPOPT reports spending 96% of its solve time in a category called “LinearSystemFactorization”.
[…]
Finally, we observe that the interior point method requires many more iterations with the full-space formulation than with the reduced-space formulation. While the theory of interior point methods’ convergence behavior with full and reduced-space formulations is not well-understood, this behavior is consistent with lower iteration counts and improved convergence reliability that have been observed for reduced-space formulations in other contexts [pacaud2022, parker2022, naik2025aggregation]. In addition to iteration counts, the times-per-iteration are significantly higher for the full-space formulation, suggesting that its bottlenecks would not be remedied by using a different optimization algorithm that is able to converge in fewer iterations.
Table 4: Solve time breakdowns for selected neural networks and formulations
Model Formulation NN. param. Platform Solve time (s) Percent of solve time (%)
Function Jacobian Hessian Solver
MNIST Full-space 5M CPU 11536 0.05 0.05 0.07 99+
MNIST Reduced-space 274M CPU 45 5 13 81 1
MNIST Reduced-space 274M CPU+GPU 3 3 5 72 19
SCOPF Full-space 4M CPU 8858 0.1 0.1 0.2 99+
SCOPF Reduced-space 592M CPU 144 6 14 80 0.2
SCOPF Reduced-space 592M CPU+GPU 3 6 19 68 8
6 Conclusion
This work demonstrates that nonlinear local optimization problems may incorporate neural networks with hundreds of millions of trained parameters, with minimal overhead, using a reduced-space formulation that exploits efficient automatic differentiation and GPU acceleration. Further research should test this formulation on neural networks with larger input and output dimensions to measure the point at which CPU-GPU data transfer becomes a bottleneck; our experiments indicate that, for our test problems, this overhead is small compared to the effort of evaluating and differentiating the neural network itself. A disadvantage of our formulation is that it is not suitable for global optimization as the non-convex neural network constraints are not represented in a format that can be communicated to any global optimization solver we are aware of. Interfacing convex under and over-estimators of neural networks (e.g., CROWN [zhang2018crown]) with global optimization solvers is another interesting area for future work. Additionally, relative performance of the full and reduced-space formulations may change in different applications. This motivates future research and development to improve the performance of the full-space formulation, which may be achieved by linear algebra decompositions that exploit the structure of the neural network’s Jacobian in the KKT matrix.
Acknowledgments and Disclosure of Funding
This work was funded by the LDRD program at Los Alamos National Laboratory under the Artimis project, the Information Science and Technology Institute, and the Center for Nonlinear Studies. LA-UR-25-29374.